QUESTION 18

Suppose Latte Larry’s successfully eliminates their rival Mocha Joe’s so that Latte Larry’s is now a coffee monopolist. Let Latte Larry’s cost function be C(q) = 2q and the inverse demand for coffee be p(q) = 20 − 0.5q.

What is the profit-maximizing price that Larry will choose?

I got P = 5 ?
Answer #1

Cost function : elg1=2q demand pla7= 20-0.5 mengind cost (me) = o = 2 Total Revenue = Pag= (20-008979 . . (ToR) TR = 20g - 0.